OpenStudy (anonymous):

1. The pet store has a treat bar that sells treats for dogs. The price for 3 lb of Carob Cookies is $8.25. 1. One customer wants to purchase 7.5 lb of Carob Cookies. How much will the store charge for the Carob Cookies? Explain

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

I am confused with this Math Question...

OpenStudy (michele_laino):

hint: what is the cost for one single lb of cookies?

OpenStudy (anonymous):

I dont know... :/

OpenStudy (anonymous):

2.??

OpenStudy (anonymous):

2.25?

OpenStudy (michele_laino):

8.25/3=...?

OpenStudy (anonymous):

2.75

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now if 1 lb costs $2.75, how much do cost 2 lb?

OpenStudy (michele_laino):

oops..how much do 2 lb cost?

OpenStudy (michele_laino):

hint: what is 2.75*2=...?

OpenStudy (anonymous):

5.5

OpenStudy (michele_laino):

ok! and how much 3 lb cost?

OpenStudy (anonymous):

8.25

OpenStudy (michele_laino):

so how much do 7.5 lb cost?

OpenStudy (anonymous):

...

OpenStudy (michele_laino):

what is 2.75*7.5=...?

OpenStudy (anonymous):

20.625

OpenStudy (michele_laino):

that's right! That is your answer!

OpenStudy (anonymous):

oh Thx!

OpenStudy (michele_laino):

Thanks!